a number when added to its half gives you 44.what is the number
Answer by Maths68(1474)

Let number = x
Half of the x = x/2
addition of x and x/2 gives 44
x+x/2= 44
Multiply by 2 both sides
2x+x=44*2
3x=88
x=88/3